OpenStudy (anonymous):

how would you solve (sqrt(x+2)) + (sqrt(3-x))

OpenStudy (anonymous):

\[\sqrt{x+2}+\sqrt{3-x}=0 ?\]

OpenStudy (anonymous):

\[\sqrt{a} ~~~\,a \ge0\]

OpenStudy (anonymous):

\[x+2\ge0 \,and \,~3-x \ge0\]

OpenStudy (anonymous):

f(x) = \[\sqrt{x+2}\] ; g(x) = \[\sqrt{3-x}\] find (f+g)(x) how would I do that?

OpenStudy (anonymous):

\[x \ge2~\, and~\,x \le3\]

OpenStudy (anonymous):

ohh ok

OpenStudy (anonymous):

(f+g)(x)=f(x)+g(x) \[(f+g)(x)=\sqrt{x+2}+\sqrt{3-x}\] this is the answer you cant do anything

OpenStudy (anonymous):

what if it was (f * g) ? could you do something then?

OpenStudy (anonymous):

since inside of the roots are different you cannot make addition

OpenStudy (anonymous):

is it multiply sign?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

if it was (f*g)(x), you can multiply the expression inside the radical signs since they have common indices.

OpenStudy (anonymous):

(f*g)(x)=f(x)*g(x) \[(f*g)(x)=(\sqrt{x+2})(\sqrt{3-x})\] \[\sqrt{(x+2)(3-x)}\]\[\sqrt{-x^2+x-6}\]

OpenStudy (anonymous):

it should be +6 not -6

OpenStudy (anonymous):

o okay that makes sense thank you

OpenStudy (anonymous):

yw